73. In a study that was performed to determine the # of vehicles using an intersection,...

73.

In a study that was performed to determine the # of vehicles using an intersection, the mean # was 375 vehicles per day and the standard deviation was 25 vehicles. Given that the distribution is symmetric and mound-shaped, what percent of the time is the # of vehicles using the intersection between 350 and 425?

A.

The correct answer is not among the choices.

B.

81.5

C.

16

D.

78.5

E.

97
Answer #1

Solution :

Given that ,

mean = = 375

standard deviation = = 25

P( 350 < x < 425) = P[(350 -375)/ 25) < (x - ) /  < (425 - 375) /25 ) ]

= P(-1 < z < 2 )

= P(z < 2) - P(z < -1)

Using standard normal table

= 0.9772 - 0.1587= 0.8186

Probability = 0.8186 = 81.86 %

Answer = 81.86 %

A . The correct answer is not among the choices.